OpenStudy (anonymous):

what is an equation of a parabola with the given vertex and focus vertex (-2 5) focus (-2 6)

OpenStudy (anonymous):

wow another conic problem 4th one tonight, and no one likes them do you know what this looks like? we need that first

OpenStudy (anonymous):

no idea

OpenStudy (anonymous):

lets plot the points at least 

OpenStudy (anonymous):

the fact that the focus is above the vertex should indicate that the parabola opens up

OpenStudy (anonymous):



OpenStudy (anonymous):

crappy picture but you get the idea i hope

OpenStudy (anonymous):

yeah i get it

OpenStudy (anonymous):

general form will be something like \[4p(y-k)=(x-h)^2\] we know the \(x\) is squared because it opens up now all you need is \(p\) and \((h,k)\)

OpenStudy (anonymous):

the vertex is \((-2,5)\) making \(h=-2,k=5\)

OpenStudy (anonymous):

so now \[4p(y-k)=(x-h)^2\]looks like \[4p(y-5)=(x+2)^2\]

OpenStudy (anonymous):

all you need is \(p\) which is the distance from the vertex to the focus the vertex is \((-2,5)\) and the focus is \((-2,6)\) which is one unit up, making \(p=1\) and the whole thing is therefore \[4(y-5)=(x+2)^2\]

OpenStudy (anonymous):

wow that was honestly really simple... i feel silly now lol. thank you

OpenStudy (anonymous):

simple once you do two or three, yes yw